OpenStudy (anonymous):

Determine the sum of the measures of the interior angles of the specified polygon. A dodecagon

OpenStudy (cwrw238):

a dodecagon has 12 sides 12 *(180 - 360/12) = 12 * 150 = 1800 degrees

OpenStudy (cwrw238):

do you understand this formula? i think theres another formula but this is the one i use

OpenStudy (anonymous):

yes, thanks i got it right

OpenStudy (cwrw238):

oh - ok - the other formula is a bit easier to use really its 180(n - 2)